OpenStudy (anonymous):

16 lb of Brand M Cinnamon was made by combining 12 lb of Indonesian cinnamon which costs $19/lb with 4 lb of Thai cinnamon which costs $11/lb. Find the cost per lb of the mixture.

OpenStudy (anonymous):

total=12x19+4x11=272 for mixture is 272/16=17 so the price is $17

OpenStudy (anonymous):

16 = 12*19 + 4*11 16lbs = 228 +44 16lbs = $272 1 lb = $17.

OpenStudy (anonymous):

wow guys... you are like machines!